### v1

# SPARTA: Advancing Sparse Attention in Spiking Neural Networks via Spike-Timing-Based Prioritization

Abstract
Current Spiking Neural Networks (SNNs) underutilize the temporal dynamics inherent in spike-based processing, relying primarily on rate coding while overlooking precise timing information that provides rich computational cues. We address this by proposing SPARTA (Spiking Priority Attention with Resource-Adaptive Temporal Allocation), which leverages heterogeneous neuron dynamics and spike-timing information to enable sparse attention mechanisms. SPARTA extracts temporal cues—including firing patterns, spike timing, and inter-spike intervals—to prioritize tokens for processing, achieving 65.4% sparsity through competitive gating. By selecting only the most salient tokens, SPARTA reduces attention complexity from $O(N^{2})$ to $O(K^{2})$, where $k\!\ll\!n$. Our approach achieves state-of-the-art accuracy on DVS-Gesture (98.78%) and competitive performance on CIFAR10-DVS (83.06%) and CIFAR-10 (95.3%), demonstrating that spike-timing utilization enables both computational efficiency and competitive accuracy.
Introduction
Deep learning with Artificial Neural Networks (ANNs) has revolutionized numerous aspects of modern society, achieving major breakthroughs in computer vision (He et al. 2016). However, the substantial energy consumption of these increasingly complex models has emerged as a critical bottleneck for practical deployment (Strubell, Ganesh, and McCallum 2019).
Spiking Neural Networks (SNNs) offer a fundamentally different paradigm by processing discrete, asynchronous spikes that mirror biological neural computation. This enables natural event-driven processing and intrinsic temporal coding, making SNNs particularly suited for neuromorphic hardware such as Intel’s Loihi (Davies et al. 2018). Yet, the discrete spike events and temporal dynamics introduce inherent challenges for training, resulting in a persistent accuracy gap compared to continuous-valued ANNs. While recent approaches have successfully improved SNN performance through ANN-inspired techniques like ANN-to-SNN conversion (Deng and Gu 2021), surrogate gradient methods (Neftci, Mostafa, and Zenke 2019), and sophisticated normalization (Zheng et al. 2021), these methods have underutilized the rich temporal dynamics inherent in spike-based processing. Most existing approaches focus primarily on achieving higher accuracy while insufficiently exploiting the precise temporal information that distinguishes SNNs from conventional ANNs, representing a key opportunity for further advancement. Consequently, this has limited their computational efficiency on neuromorphic hardware (Orchard, Delbruck, and Davies 2021; Bellec et al. 2018). SNNs inherently excel in spatio-temporal coding, leveraging precise spike timing to efficiently encode complex temporal patterns, a capability ANNs do not naturally replicate (Eshraghian et al. 2023). However, current research predominantly focuses on matching ANN performance metrics while insufficiently addressing the unique opportunities that spike timing provides for both computational efficiency and attention mechanisms. This raises a critical question: How can we more effectively integrate the temporal coding capabilities inherent to spike dynamics into attention mechanisms to enhance both efficiency and performance in SNNs?
Inspired by neuroscientific evidence from insights into temporal dynamics underlying selective attention (Singer 1999), we propose SPARTA (Spiking Priority Attention with Resource-Adaptive Temporal Allocation). SPARTA incorporates heterogeneous neurons that mimic the rich diversity of response properties found in cortical neuron populations, enabling the network to capture complex temporal features across multiple timescales. It then leverages a Spatio-Temporal Encoding Network (STEN) to construct a multi-scale feature representation that explicitly preserves this critical spike timing information. Finally, these features guide a Priority-Aware Sparse Temporal Attention, which dynamically allocates computational resources only to the most salient tokens, avoiding the quadratic complexity of standard attention while maintaining processing efficiency.
SPARTA’s token selection mechanism is based on three biologically-inspired observations: (1) important stimuli tend to fire earlier (Foffani, Chapin, and Moxon 2009), (2) important stimuli tend to fire with shorter intervals between spikes (Oswald et al. 2007), and (3) important stimuli tend to fire more frequently (Gerstner et al. 1997). These observations guide our multi-scale feature extraction approach. We interpret stimuli as tokens, enabling biologically-inspired selection and localized competitive gating for sparse attention, while preserving event-driven sparsity critical to neuromorphic hardware.
Our contributions are:
  • Heterogeneous Initialized Leaky Integrate-and Fire (HI-LIF) neuron that introduces learnable, channel-wise temporal diversity to expand the network’s processing bandwidth.
  • A novel priority-aware sparse temporal attention mechanism, guided by biologically-inspired cues (e.g., firing rate, spike timing), for efficient, saliency-based computation.
  • The SPARTA framework, which achieves competitive performance demonstrating that integrating biologically-inspired temporal cues can enhance both efficiency and performance in SNNs.
Background and Motivation
LIF Neuron Models and Temporal Coding
The computational core of Spiking Neural Networks (SNNs) is the Leaky Integrate-and-Fire (LIF) neuron, whose dynamics are governed by a membrane time constant ($\tau$) and a firing threshold ($v_{th}$), as described in Equations 1 and 2.
$u^{(t+1)}=u^{(t)}\left(1-\frac{1}{\tau}\right)+x^{(t)}$ (1)
$s^{(t)}=\Theta\left(u^{(t)}-v_{th}\right)$ (2)
However, a critical limitation arises in how these models are conventionally applied: SNNs typically employ uniform parameters shared across all spatial channels. This simplification, while computationally convenient, starkly contrasts with biological reality, where cortical neurons exhibit remarkable heterogeneity in their temporal properties (Mason et al. 2022; Eyal et al. 2023). This imposed homogeneity creates a significant bottleneck, limiting the network’s temporal coding capacity and its ability to process complex information across multiple timescales (Perez-Nieves et al. 2021). Our work directly confronts this limitation, proposing a neuron model inspired by this biological diversity to unlock a richer temporal processing bandwidth.
Human Cognition and Temporal Attention
Human visual attention leverages temporal dynamics for rapid pattern recognition. The flashed-face distortion effect demonstrates temporal sensitivity: faces presented in rapid succession appear perceptually distorted, consistent with competitive normalization within brief presentation windows (Tangen, Murphy, and Thompson 2011). Structured visual search tasks illustrate temporal attention mechanisms where systematic scanning operates under specific temporal constraints (Wolfe 1994; Chun and Potter 1995). Cognitive integration depends on maintaining partial cues within critical time windows, beyond which integration success declines (Chun and Potter 1995; Di Lollo 1977). While these cognitive phenomena provide a high-level, conceptual foundation for our approach, the core computational mechanisms of SPARTA are grounded in established models from computational neuroscience.
Figure 1: Crossword puzzle analogy illustrating temporal integration and decay of visual cues underlying SPARTA’s selective temporal attention.
Figure 1 illustrates a step-by-step temporal attention process analogous to our problem:
  1. 1.
    Global scan: Continuous monitoring detects “FAT” but dismisses it, as “FAT” is not a valid psychological word.
  2. 2.
    Rapid detection: Focused attention shifts to fragment “IGUE” within the relevant time interval.
  3. 3.
    Temporal integration: The brain associates “FAT” and “IGUE” to form “FATIGUE” (a valid psychological term).
  4. 4.
    Interval dependency: Longer delays cause the memory of “FAT” to decay, reducing the likelihood of integration. The user will be unable to connect “IGUE” with “FAT”.
Such observations suggest how temporal windows may affect the selection and integration of sensory cues, providing a conceptual foundation for SPARTA’s spike-timing based attention mechanism. This temporal integration process parallels spike-based neural computation, where information binding occurs through precise timing relationships rather than simple accumulation. Unlike conventional frame-based approaches that process each temporal snapshot independently, spiking networks naturally maintain temporal context through membrane dynamics, enabling information integration across biologically plausible time windows.
Building on this principle, SPARTA leverages spike timing to preserve the temporal dynamics essential for efficient attention allocation, drawing inspiration from both cognitive processes and neuromorphic computation principles.
Event-based Vision
Event-based cameras (e.g., Dynamic Vision Sensors; DVS) emit polarity events only when local log-intensity changes occur, producing temporally precise and highly sparse streams that interface cleanly with spike-based SNN computation. (Lichtsteiner, Posch, and Delbruck 2008; Gallego et al. 2020). These sensors capture key aspects of change-driven encoding found in biological vision systems (Delbruck and Lichtsteiner 2014; Gollisch and Meister 2010). Learning directly from native event streams is practical at scale; accumulating events into dense frames discards fine-grained timing info and can substantially increase memory and computational demands (Amir et al. 2017; Gallego et al. 2020). Neuromorphic processors (e.g., Loihi 2) achieve substantial energy savings by scheduling work only on active addresses—idle pixels and silent neuron populations incur negligible cost—making event-driven sparsity a primary efficiency lever (Davies et al. 2018; Orchard, Delbruck, and Davies 2021).
Related Works
Performance-Driven SNNs focus on accuracy through ANN-inspired techniques, primarily emphasizing rate-based coding. ANN-to-SNN conversion methods achieve high performance through weight scaling and threshold adjustment (Bu et al. 2022; Han, Srinivasan, and Roy 2020; Kim et al. 2022). Hybrid training schemes apply gradient-based optimization to spike networks (Li et al. 2021). Large-scale adaptations port state-of-the-art models to spiking paradigms (Yao et al. 2023; Zheng et al. 2023; Wang et al. 2023), demonstrating compatibility with cutting-edge AI while treating spikes primarily as discrete rate codes rather than exploiting their temporal coding capabilities.
Biologically-Inspired Architectures emphasize neuromorphic principles and biological similarity in design. Local learning rules employ STDP and Hebbian mechanisms for biological fidelity (Song, Miller, and Abbott 2000; Diehl and Cook 2015), offering hardware compatibility while optimizing for specific learning paradigms that prioritize biological authenticity. Neuromorphic hardware designs achieve energy efficiency and real-time processing through platform-specific optimizations for systems like Loihi (Davies et al. 2018) and SpiNNaker (Furber et al. 2014), demonstrating effective integration of neuromorphic principles with practical hardware constraints. Brain-circuit architectures directly emulate neural circuits through cortical column simulations (Hawkins and Ahmad 2016), achieving high biological similarity while specializing in applications where biological fidelity is the primary design criterion.
Temporal Dynamics & Sparse Processing focuses on preserving and leveraging spike timing information to optimize computational efficiency and accuracy. Methods like optimized spiking neurons achieve high accuracy through precise timing codes (Stockl and Maass 2021), while sparse processing frameworks reduce energy consumption without sacrificing network performance (Yin, Corradi, and Bohté 2023). Multi-scale encoding approaches expand receptive fields across resolution levels (Dampfhoffer et al. 2021), and progressive learning methods enable deep networks to process complex patterns through sparse representations (Wu et al. 2021). Recent empirical analysis demonstrates temporal information dynamics in SNNs, showing natural concentration in earlier timesteps during training (Kim et al. 2023). Spatio-temporal attention mechanisms effectively integrate temporal dependencies without additional computational overhead (Lee et al. 2025). While these approaches demonstrate significant progress in leveraging temporal information, there remains opportunity to further integrate biological principles with sparse attention mechanisms to achieve balance between temporal coding and computational efficiency.
Methodology
HI-LIF: Heterogeneous Temporal Dynamics. Real cortical neurons display substantial variability in both membrane time constants ($\tau$) and firing thresholds ($v_{th}$), enabling concurrent processing of fast transients and long‑range context (Mason et al. 2022; Eyal et al. 2023). To endow SNNs with richer temporal bandwidth, we propose the Heterogeneous Initialized Leaky Integrate‑and‑Fire (HI‑LIF) neuron, which samples $\tau$ and $v_{th}$ per channel from learnable normal priors. This channel‑wise diversity expands the temporal receptive field while preserving event‑driven sparsity.
Algorithm 1 HI-LIF: Heterogeneous Temporal Dynamics
Input: Input current $x^{(t)}$ with shape $N\times C\times H\times W$
Parameters: $\mu_{\tau}$, $\sigma_{\tau}$ (tau distribution), $\mu_{v_{th}}$, $\sigma_{v_{th}}$ (threshold distribution), reset_mode (hard or soft)
Output: Spike output $s^{(t)}$ and updated membrane potential $v^{(t+1)}$
1:Initialization:
2:for each channel $c\in\{1,2,\ldots,C\}$ do
3:  $\tau_{init}^{(c)}\sim\mathcal{N}(\mu_{\tau},\sigma_{\tau}^{2})$, $\tau_{init}^{(c)}=\max(\tau_{init}^{(c)},1.01)$
4:  $w^{(c)}=-\log(\tau_{init}^{(c)}-1.0)$, $v_{th}^{(c)}\sim\mathcal{N}(\mu_{v_{th}},\sigma_{v_{th}}^{2})$
5:end for
6:Forward Pass:
7:for each channel $c$ do
8:  $\tau_{inv}^{(c)}=\sigma(w^{(c)})$
9:  $v^{(t+1,c)}=v^{(t,c)}+(x^{(t,c)}-v^{(t,c)})\times\tau_{inv}^{(c)}$Neuronal charge}
10:  $s^{(t,c)}=\Theta(v^{(t+1,c)}-v_{th}^{(c)})$Neuronal fire}
11:  if $s^{(t,c)}=1$ then
12:   if reset_mode = hard then
13:    $v^{(t+1,c)}=v_{reset}$Hard reset to fixed value}
14:   else
15:    $v^{(t+1,c)}=v^{(t+1,c)}-v_{th}^{(c)}$Soft reset with channel-wise threshold subtraction}
16:   end if
17:  end if
18:end for
Figure 2: SPARTA architecture overview. Input events are processed through STEN to generate SpikeInfo with rich temporal cues. Following multi-scale processing by MSP, the STSG module modulates token importance, after which the SC selects the top-K salient tokens for efficient sparse attention classification. A feedback controller (dashed lines) dynamically adjusts STEN parameters. Solid arrows show the main data flow.
[…]
Multi-Scale Spike Encoding & Feature Extraction
As mentioned in introduction, SPARTA’s token selection pipeline ranks patch saliency using three biologically driven cues—high firing rate, early first‑spike, and short inter‑spike interval.
Spatio-Temporal Encoding Network (STEN). STEN implements cascading downsampling that preserves temporal dynamics through HI-LIF neurons. It processes spike features through three parallel branches: (i) 1×1 convolution for fine-grained details, (ii) 3×3 convolution with HI-LIF for additional dynamics, and (iii) adaptive pooling for global context. The resulting multi-scale features are concatenated and refined by a timing-aware attention mechanism, which assigns higher weights to rich temporal activity. In parallel, STEN derives three complementary timing metrics—first-spike timing $T_{\mathrm{first}}$ for rapid event detection, inter-spike intervals $T_{\mathrm{interval}}$ for continuity, and burst firing patterns $T_{\mathrm{burst}}$ for salience estimation, forming a comprehensive temporal representation for downstream processing.
Multi-Scale Processing (MSP). The second stage applies bias-based attention mechanisms that weight different temporal characteristics. This approach is grounded in established temporal coding models from computational neuroscience. Specifically, we model the decaying importance of spike latency using an exponential function, a common and effective method in temporal plasticity models (Szatmáry and Izhikevich 2010; Zhang et al. 2024; Diehl and Cook 2015). The weights are defined as:
$w^{(s)}_{\mathrm{timing}}=\exp(-\alpha\cdot T^{(s)}_{\mathrm{first}})$ (3)
$w^{(s)}_{\mathrm{interval}}=\exp(-\beta\cdot T^{(s)}_{\mathrm{interval}})$ (4)
$w^{(s)}_{\mathrm{combined}}=w^{(s)}_{\mathrm{timing}}\odot w^{(s)}_{\mathrm{interval}}\odot\sigma(\gamma\cdot F^{(s)}_{\mathrm{rate}})$ (5)
Here, the exponential formulation for $w^{(s)}_{\mathrm{timing}}$ directly models the principle that earlier spikes carry greater informational value, a key aspect of first-spike latency codes (Foffani, Chapin, and Moxon 2009; Guo et al. 2021). The factors $\alpha,\beta,\gamma$ are learnable scaling parameters that allow the network to adaptively balance these complementary temporal cues. The resulting temporal bias is incorporated into multi-head attention through attention masking to emphasize temporally salient regions.
Patch Grouping. Aggregates multi-scale spike features and adjusts the token count to a fixed size without zero-padding by selecting or duplicating tokens based on their importance (e.g., firing rate). This preserves meaningful spike information while ensuring compatibility with downstream sparse attention modules.
Sparse Token Processing & Attention
[…]
Spike Token Selection & Gating (STSG). STSG implements lateral inhibition mechanisms that integrate three attention mechanisms: MSP features, spatial competition through center-surround inhibition kernels, and temporal priority information. Unlike fixed sparsity ratios, STSG employs a learned predictor that adapts to temporal characteristics:
$\mathbf{f}_{\mathrm{input}}=\left[\begin{array}[]{c}\mathrm{mean}(F_{\mathrm{rate}})\\ \mathrm{std}(T_{\mathrm{timing}})\\ \mathrm{mean}(T_{\mathrm{interval}})\end{array}\right]$ (6)
$\rho_{\mathrm{dynamic}}=\sigma\!\left(\mathrm{MLP}\!\left(\mathbf{f}_{\mathrm{input}}\right)\right)$ (7)
The three scoring mechanisms are fused through a learned attention network:
$\mathbf{s}_{\mathrm{combined}}=\mathrm{MLP}_{\mathrm{fusion}}([\mathbf{s}_{\mathrm{spatial}},\mathbf{s}_{\mathrm{MSP}},\mathbf{s}_{\mathrm{temporal}}])$ (8)
The dynamic K value is computed based on predicted sparsity ratio with a minimum threshold for stable processing, and top-K tokens are selected based on fused attention scores. The lateral inhibition mechanism applies differential processing where selected tokens receive enhancement while non-selected tokens undergo suppression:
$\mathbf{f}_{\mathrm{processed}}=\mathbf{f}\odot\mathbf{M}_{\mathrm{topK}}\times\alpha+\mathbf{f}\odot(1-\mathbf{M}_{\mathrm{topK}})\times\beta$ (9)
where $\mathbf{M}_{\mathrm{topK}}$ is the binary selection mask, and $\alpha$, $\beta$ are learned enhancement and suppression factors respectively.
[…]
Feedback Controller. To maintain stable network activity and prevent saturation, SPARTA incorporates a feedback controller that dynamically adjusts the firing thresholds of the HI-LIF neurons based on an exponential moving average of activity from the sparse attention layers. This adaptive mechanism ensures the network maintains optimal firing rates without manual tuning. The detailed implementation is provided in Appendix.
Experiments
To validate the effectiveness of our approach, we conduct experiments on neuromorphic datasets (DVS Gesture (Amir et al. 2017), CIFAR10-DVS (Li et al. 2017)) and conventional RGB datasets (CIFAR-10 (Krizhevsky and Hinton 2009), CIFAR-100 (Krizhevsky and Hinton 2009)). Our evaluation focuses on analyzing SPARTA’s overall performance and temporal variance of model.
Experimental Setup. Experiments were performed using PyTorch with AdamW, CrossEntropy Loss (lr 1e-4, cosine schedule); neuromorphic tasks ran 300 epochs, RGB tasks 500 epochs. Results are the mean of three seeds.
Figure 3: Visualization of gesture samples.(a) Input frame from DVS-Gesture dataset (top: arm-roll, bottom: clap gestures); (b) Firing rate map (red to yellow: low to high rates); (c) Attention weights (purple to yellow: low to high); (d) Top-K selection mask (white: selected tokens, black: filtered tokens); (e) Spike count variance across temporal dimension. The arm-roll gesture shows uniform spike counts over time, while clap exhibits concentrated spike bursts.
Ablation Study
We conduct systematic ablation studies to understand each component’s contribution to SPARTA’s effectiveness, beginning with temporal resolution analysis and progressing through individual component evaluations.
Study on Temporal Resolution.
We first analyze SPARTA’s performance across different time step configurations to understand its temporal processing characteristics and establish optimal operating conditions.
T Accuracy (%) Variance
DVS-Gesture CIFAR10-DVS Timing / Interval
4 88.89 (-10.01%) 78.2(-6.07%) 0.287 / 1.81
8 92.74 (-6.04%) 78.7(-5.45%) 0.288 / 1.93
12 94.73 (-4.05%) 81.90(-1.45%) 0.280 / 1.92
16 98.46 (-0.32%) 83.06 0.288 / 1.71
20 98.78 82.87(-0.24%) 0.292 / 1.82
32 94.30 (-4.54%) 79.45(-4.51%) 0.287 / 1.91
Table 1: Temporal performance at different timesteps ($T$). The ‘Variance‘ column reports the spatial variance of first-spike timings and inter-spike intervals, respectively, computed across all output tokens and averaged over the entire test set.
Results show optimal performance at T=20 for DVS-Gesture (98.78%) and T=16 for CIFAR10-DVS (83.06%). Both datasets exhibit performance degradation at T=32 (-4.54% and -4.51%), indicating that excessive temporal windows introduce noise and reduce accuracy.
HI-LIF Heterogeneity Analysis.
We analyze the impact of channel-wise heterogeneity on temporal encoding efficacy by varying the standard deviations of the membrane time constant ($\tau$) and firing threshold ($v_{th}$).
Configuration HI-LIF Parameters Accuracy
$\tau_{\sigma}$ $v_{th,\sigma}$ (%)
Homogeneous 0.0 0.0 96.53
Low Tau Diversity 0.2 0.0 97.25
High Tau Diversity 0.5 0.0 94.36
Low Threshold Diversity 0.0 0.1 96.27
High Threshold Diversity 0.0 0.3 92.43
Combined Diversity 0.5 0.3 89.83
Adjusted (Combined) 0.3 0.2 98.46
Table 2: Impact of individual HI-LIF parameter diversity on accuracy(DVS-Gesture) ($\tau_{\mu}=2.0$, $v_{th,\mu}=1.0$, T=16).
Results show that adjusted combined heterogeneity achieves the highest accuracy, demonstrating the benefit of balanced diversity, while excessive diversity leads to performance degradation due to instability.
Analysis of Sparsity Policies.
We benchmark our dynamic sparsity policy against fixed-sparsity baselines on the DVS-Gesture dataset ($N=256$), evaluating the trade-off between accuracy and computational cost.
Sparsity Policy Sparsity (%) Acc. (%) FLOPs (G)
Dynamic (Ours) 65.4 (Adaptive) 98.78 1.23
Fixed (K=192) 25.0 98.30 1.24
Fixed (K=128) 50.0 92.50 1.21
Fixed (K=64) 75.0 78.03 1.19
Table 3: DVS-Gesture accuracy and computational cost for dynamic vs. fixed sparsity policies (T=20). For fixed policies, K denotes the number of tokens resulting from the specified sparsity, calculated as $N\times(1-\mathrm{Sparsity}/100)$.
Our dynamic policy selector achieves 65.4% average sparsity while maintaining 98.78% accuracy, outperforming all fixed baselines. As shown in Table 3, our adaptive policy avoids the trade-off seen in fixed-sparsity approaches: a low fixed sparsity (high K) achieves high accuracy at a higher computational cost, while a high fixed sparsity (low K) minimizes FLOPs but suffers from a significant drop in accuracy. The learnable sparsity predictor optimally balances these objectives by adapting the number of selected tokens ($K$) to input complexity.
MSP Temporal Weighting Ablation.
We systematically ablate the temporal weighting parameters in MSP to understand the contribution of each biological cue.
Configuration T DVS-Gesture CIFAR10-DVS
Full MSP ($\alpha$,$\beta$,$\gamma$) 16 98.46 83.06
w/o $\alpha$ (timing) 16 96.94 (-1.52) 79.80 (-3.26)
w/o $\beta$ (interval) 16 95.56 (-2.90) 80.35 (-2.71)
w/o $\gamma$ (firing rate) 16 92.26 (-6.2) 77.23 (-5.83)
w/o $\alpha$,$\beta$ 16 94.61 (-3.85) 78.28 (-4.78)
w/o $\alpha$,$\gamma$ 16 89.16 (-9.30) 75.26(-7.80)
w/o $\beta$,$\gamma$ 16 91.38 (-7.08) 74.50 (-8.56)
w/o $\alpha$,$\beta$,$\gamma$ 16 85.0 (-13.46) 72.80 (-10.26)
Table 4: Ablation study on MSP’s temporal weighting cues (Accuracy %). $\alpha$, $\beta$, and $\gamma$ correspond to the weights for first-spike timing, inter-spike interval, and firing rate, respectively.
Table 4 reveals a clear hierarchy of importance among the temporal cues. The firing rate weight ($\gamma$) is consistently the most critical single factor, as its removal causes the largest individual performance drop on both DVS-Gesture (-6.2%) and CIFAR10-DVS (-5.83%). While timing ($\alpha$) and interval ($\beta$) cues also contribute significantly, the most substantial degradation occurs when all three are removed entirely. This confirms that these weights work in a complementary manner to effectively guide the attention mechanism.
Method Params (M) DVS-Gesture CIFAR10-DVS CIFAR-10 CIFAR-100
T Acc T Acc T Acc T Acc
SEW-ResNet (Fang et al. 2021) 60.2 16 89.06 16 67.20 - - 4 75.93
GLIF+ResNet (Yao et al. 2022) 11.2 - - 16 78.10 4 94.67 4 77.37
Spikformer (Zhou et al. 2023) 9.32 16 95.49 16 80.60 4 95.19 4 77.86
SpikingResFormer (Shi, Hao, and Yu 2024) 17.25 16 91.67 10 84.80 4 97.40 4 85.98
QKFormer-S/L (Zhou et al. 2024) 1.5/6.74 16 98.60 16 84.00 4 96.18 4 81.15
SGLFormer (Zhang et al. 2024) 8.9 16 97.20 10 82.90 4 96.76 4 82.26
Event-Vivid (Li et al. 2024) 48.2 20 98.80 20 92.50 - - - -
SMA-AZO-VGG (Shan et al. 2025) - 16 98.60 10 84.00 - - - -
SPARTA(Ours) 13.8 20 98.78 16 83.06 4 95.3 4 78.1
Table 5: State-of-the-art comparison on neuromorphic and RGB datasets. T denotes the timesteps; Acc (%) indicates classification accuracy; Params refers to the number of model parameters. A dash (–) indicates values not reported in the original paper.
Temporal Variance Analysis.
To further validate SPARTA’s temporal encoding capabilities, we analyze its spike timing variance against other SNN architectures (Perez-Nieves et al. 2021; Bellec et al. 2018). Specifically, we measure the spatial variance of both first-spike timings and inter-spike intervals across all output tokens, then average these values over the entire DVS-Gesture test set. This analysis follows established methodologies for evaluating temporal diversity in spiking networks (Stockl and Maass 2021; Mason et al. 2022).
Figure 4: Temporal variance comparison across different SNN architectures on the DVS-Gesture dataset. The y-axis represents the spatial variance of spike timings and intervals, averaged over the test set.
As shown in Figure 4, SPARTA maintains higher temporal diversity than the compared models. This confirms that the heterogeneous parameters of our HI-LIF neurons enable the network to learn richer and more diverse temporal representations, which is a key factor in its strong performance.
Comparison with State-of-the-Art
We conduct comprehensive experiments to compare our proposed SPARTA method with recent state-of-the-art SNN models across four benchmark datasets. Table 5 presents the performance comparison on both neuromorphic datasets (DVS-Gesture, CIFAR10-DVS) and static RGB datasets(CIFAR-10, CIFAR-100).
Conclusion
We present SPARTA, a biologically-inspired Spiking Neural Network (SNN) framework designed to bridge the gap between neuromorphic fidelity and competitive performance. SPARTA leverages Heterogeneous Initialized Leaky Integrate-and-Fire (HI-LIF) neurons to enhance temporal processing diversity and employs a Spatio-Temporal Encoding Network (STEN) to extract and preserve critical temporal information. The framework implements a two-stage sparse attention mechanism: first, the Spike Token Selection &Gating (STSG) performs temporal modulation of input tokens by weighting them according to biologically-motivated cues, followed by the Sparse Classifier (SC) that selects the top-K most salient modulated tokens and performs final classification with $O(K^{2})$ computational complexity. This sequential processing pipeline ensures that biological temporal dynamics are preserved while achieving computational efficiency. Our experimental results on DVS-Gesture (98.78%) and CIFAR10-DVS (83.06%) validate the core hypothesis that biological principles can coexist with high accuracy in large-scale SNNs. However, we acknowledge several limitations: the sparse attention mechanism may lead to information loss when critical information is distributed across many tokens (e.g., fine-grained textures) or when important temporal patterns occur in tokens that are filtered out during top-K selection, and its performance has only been validated on classification tasks. Future work will focus on extending SPARTA to multi-modal data, and deploying the framework on neuromorphic hardware to verify its real-world efficiency.